OpenStudy (anonymous):

A cylinder has radius HC and height h. What is the ratio of the volume of the paraboloid to the volume of the cylinder?

OpenStudy (anonymous):

 this the diagram

OpenStudy (raden):

the paraboloid of y=x^2 have rounded x-axes or y-axes ?

OpenStudy (anonymous):

y-axis and the volume for the parabloid is pih^2/2

OpenStudy (raden):

yeps... the volume of parabolid : Vp = pi * int (sqrt(y))^2 dy [0,h] = pi * int (y) dy [0,h] = pi * 1/2* y^2 [0,h] = pi*1/2*h^2 Next, to find the volume of cylinder, first we must determine the intersection between line y=h and y=x^2 x^2=h x=+-sqrt(h) (nevermind the negative sign, because radius impossible -) so, the radius of cylinder r=sqrt(h) just doing by manual, :p Vc = pi*r^2*t = pi*(sqrt(h))^2*h =pi*h^2 now u can get ratio of both : Vp : Vc = ....

OpenStudy (raden):